[A] Consider the inner product space obtained by equipping ?[0,2] with the inner product given below:
〈?(?),?(?)〉 = ∫ ?(?)?(?)?? 2 0


Determine the value of each of the following (simplifying where possible; no decimals). You do not have to show the steps of calculating the integrals, but must at least write the integrals used in your calculations.   

(A.1) 〈?,1〉   




(A.2) ‖ ? − 1 ‖

(A.3) ?(??,?? + 10), i.e. the distance between ?? and ?? + 10 .   







(A.4) Determine the value of ? so that the functions ?(?) = ? and ?(?) = ?? + 2 are  orthogonal relative to the given inner product.

(A.5) Consider the subspace ℙ1 of ?[0,2] consisting of all polynomials of degree 1 or less. Use the Gram-Schmidt process to convert its standard basis ? = {1,?} to an orthonormal basis.
Answer #1

A.1. By definition,

\langle t,1\rangle = \int_0^2 t\cdot 1 \,\text dt
2 = ſ tdt
=\left[ \frac{t^2}{2} \right ]_0^2
= 2 .

A.2

||t – 1|| = Vit - 1,t - 1)
=\sqrt{ \int_0^2 (t-1)\cdot (t-1) \,\text dt }
=\sqrt{ \int_0^2 (t-1)^2 \,\text dt }
=\sqrt{ \int_{-1}^{1} \tau^2 \,\text d\tau }
=\sqrt{\left[ \frac{\tau^3}{3} \right ]_{-1}^1 }
=\sqrt{\frac 23 } .

A.3

det, et + 10) = || (e + 10) -(e)
= \| 10\|
=\sqrt{\int_0^2 10 \cdot 10 \,\text dt}
= 10 \sqrt 2 .

A.4

We want to find a such that

0 = \langle t,at+2\rangle
=\int_0^2 t\cdot(at+2)\,\text dt
=\int_0^2 at^2 +2t\,\text dt
= \left[ a\frac{t^3}3 +t^2\right ]_0^2
=\frac{8}{3}a + 4 .

Therefore, a = -\frac 32 .

A.5 We define u_1 = 1 and u_2 = t . We want to convert S = \{u_1,u_2\} into a orthonormal basis \{v_1,v_2\} . First we normalize u_1 to get v_1 ,

v_1 = \frac{u_1}{\|u_1\|}
= \frac{1}{\|1\|}
= \frac{1}{\sqrt{\int_0^2 1 \,\text dt}}
=\frac{1}{\sqrt{2}}
=\frac{\sqrt{2}}2

Now we define

V_2 = v_2 - \text{proj}_{v_1} (u_2)
= u_2 - \langle u_2, v_1\rangle v_1
= t -\left( \int_0^2 \frac{\sqrt{2}}2 t \,\text dt\right)\frac{\sqrt{2}}2
= t - \frac 12 \int_0^2 t\,\text dt
= t - \frac 12\left[ \frac{t^2}{2} \right]_0^2
= t -1 .

We just need to normalize V_1 ,

v_1 = \frac{V_1}{\|V_1\|}
= \frac{t-1}{\|t-1\|}

From A.2 we know that \|t-1\| = \sqrt{\frac23} . Therefore,

v_1 = \sqrt{\frac 32} (t-1).

The orthonormal basis is then \left\{\frac{\sqrt 2}{2}, \sqrt{\frac 32} (t-1)\right\} .